OpenStudy (anonymous):

What are the solutions for x in the proportion 2x+3 = 2x ---- ---- x 2 What are the solutions for x in the proportion 2x+3 = 2x ---- ---- x 2 @Mathematics

jhonyy9 (jhonyy9):

4x+6=2x2 2x2-4x-6=0 x2-2x-3=0

OpenStudy (anonymous):

x = 3 or, -1

jhonyy9 (jhonyy9):

(x-3)(x+1) x=3 x=-1

OpenStudy (anonymous):

x^2 -3x+x-3=0 x(x-3) +1(x-3) = 0 (x-3)(x+1) = 0 therefore x = 3 or x = -1

OpenStudy (anonymous):

explain the two ways to simplify \[\sqrt{320}\] and provide the simplified form.

OpenStudy (aroub):

1.cross multiply 2.equation:2(2x+3)=2x^2 3.slove:4x+12=2x^2 4x+12-2x^2=0 (x-3)(x+1) x=3 & x=-1 :D

OpenStudy (aroub):

that's 320?

OpenStudy (anonymous):

\[8\sqrt{5}\]

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

320 = 16 * 20 = 4^2 * 2^2 * 5 sqrt{320} = 8Sqrt{5}

OpenStudy (anonymous):

Explain

OpenStudy (anonymous):

very simple. result: \[2x ^{2}-4x-6\] values: x=3 and x=-1